OpenStudy (anonymous):

Taylor Polynomials Let $f \in \mathscr{C}^{(m)}(E),$ where $E$ is an open subset of $R^{n}$. Fix \textbf{a} $\in E$, and suppose \textbf{x} $\in R^{n}$ is so close to \textbf{0} that the points \begin{equation*} \textbf{p}(t)=\textbf{a}+t\textbf{x} \end{equation*} lie in $E$ whenever $0 \leq t \leq 1$. Define \begin{equation*} h(t)=f(\textbf{p}(t)) \end{equation*} for all t $\in R^{1}$ for which \textbf{p}(t) $\in E$. (a) For $1 \leq k \leq m$, show (by repeated application of the chain rule) that \begin{equation*} h^{(k)}(t)=\sum(D_{i_{1} \dots i_{k}}f)(\textbf{p}(t))x_{i_{1} \dots

OpenStudy (anonymous):

Let $f \in \mathscr{C}^{(m)}(E),$ where $E$ is an open subset of $R^{n}$. Fix \textbf{a} $\in E$, and suppose \textbf{x} $\in R^{n}$ is so close to \textbf{0} that the points \begin{equation*} \textbf{p}(t)=\textbf{a}+t\textbf{x} \end{equation*} lie in $E$ whenever $0 \leq t \leq 1$. Define \begin{equation*} h(t)=f(\textbf{p}(t)) \end{equation*} for all t $\in R^{1}$ for which \textbf{p}(t) $\in E$. (a) For $1 \leq k \leq m$, show (by repeated application of the chain rule) that \begin{equation*} h^{(k)}(t)=\sum(D_{i_{1} \dots i_{k}}f)(\textbf{p}(t))x_{i_{1} \dots}x_{i_{k}}. \end{equation*} The sum extends over all ordered $k$-tuples $(i_{1, \dots ,}i_{k}$ in which each $i_{j}$ is one of the integers $1,\dots , n$.\\ ($b$) By Taylor's theorem (5.15), \begin{equation*} h(1)=\sum_{k=0}^{m-1}\frac{h^{(k)}(0)}{k!}+\frac{h^{(m)}(t)}{m!} \end{equation*} for some $t \in (0,1)$. Use this to prove Taylor's theorem in $n$ variables by showing that the formula \begin{equation*} f(\textbf{a}+\textbf{x})=\sum_{k=0}^{m-1}\frac{1}{k!}\sum(D_{i_{1}...i_{k}}f)(\textbf{a})x_{i_{1}...}x_{i_{k}}+r(\textbf{x}) \end{equation*} represents $f$(\textbf{a}+\textbf{x}) as the sum of its so-called ``Taylor polynomial of degree $m-1$", plus a remainder that satisfies \begin{equation*} \lim_{x \rightarrow 0}\frac{r(\textbf{x})}{|\textbf{x}|^{m-1}}=0. \end{equation*} Each of the inner sums extends over all ordered $k$-tuples $(i_{i,...,}i_{k})$, as in part $(a)$; as usual, the zero-order derivative of $f$ is simply $f$, so that the constant term of the Taylor polynomial of $f$ at \textbf{a} is $f$(\textbf{a}).\\ $(c)$ Exercise 29 shows that repetition occurs in the Taylor polynomial as written in part $(b)$. For instance, $D_{113}$ occurs three times, as $D_{113}$, $D_{131}$, $D_{311}$. The sum of the correspondin gthree terms can be written in the form \begin{equation*} 3(D_{1}^{2}D_{3}f)(\textbf{a})x_{1}^{2}x_{3}. \end{equation*} Prove (by calculating how often each derivative occurs) that the Taylor polynomial in $(b)$ can be written in the form \begin{equation*} \sum\frac{(D_{1}^{s_{1}} \dots D_{n}^{s_{n}}f)(\textbf{a})}{s_{1}! \dots s_{n}!}x_{1}^{s_{1}} \dots x_{n}^{s_{n}} \end{equation*} Here the summation extends over all ordered $n$-tuples $(s_{1,...,}s_{n})$ such that each $s_{i}$ is a nonnegative integer, and $s_{1}+ \dots + s_{n} \leq m-1$.